OpenStudy (anonymous):

x^4 + x^3 + x^2 + x + 1 How do i factorize this so I can find the roots of this polynomial? Thanks for checking it out!

hero (hero):

You can't factorize that, sorry

hero (hero):

Fiddle is probably trying to figure out a way to solve this :P

OpenStudy (anonymous):

x^4+x^3+x^2+x+1 =x^2(x^2+x)+(x^2+x)+1 =(x^2+1)(x^2+x)+1 Looking for roots: (x^2+1)(x^2+x)+1=0 (x^2+1)(x^2+x)=-1 x^2 >=0 --> x^2+1 is always positive x^2 >= x --> x^2+x >= 0 --> x^2+x is never negative (always positive number ) * (never negative number) = -1 is a FALSE statement so no real roots.

hero (hero):

*Sigh*

OpenStudy (nikvist):

\[x^4+x^3+x^2+x+1=0\quad,\quad x\neq 1\]\[x^4+x^3+x^2+x+1=0\quad/\cdot(x-1)\]\[x^5-1=0\]\[x^5=1=e^{2\pi k}\quad,\quad k=1,2,3,4\]\[x=e^{2\pi k/5}\quad,\quad k=1,2,3,4\]

OpenStudy (anonymous):

nikvist that was brilliant! thank you so much